OpenStudy (anonymous):

the perimeter of a square EFGH is 48 inches and EF=6x-42 what is the value of x?

OpenStudy (anonymous):

Can you solve for \( x \) in \( 4 (6x-42)=48 \) ?

OpenStudy (anonymous):

(24x-42)-48?

OpenStudy (anonymous):

Yes but clever one is \( 6x-42=12 \Rightarrow x= \frac{54}6 \)

OpenStudy (anonymous):

so x equals 54?

OpenStudy (anonymous):

no x=54/6=9

OpenStudy (anonymous):

\( x-42=12 \Rightarrow x= \frac{54}6 \Rightarrow x = 9 \)